how do you solve function notation if h(x)=x to the second power-2x+1

Answer by Alan3354(69443)

how do you solve function notation if h(x)=x to the second power-2x+1
-------
If you mean find the zeroes of x^2 - 2x + 1:
x^2 - 2x + 1 = 0
(x-1)^2 = 0
x = 1 (multiplicity 2)